OpenStudy (anonymous):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold. Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work. Describe how you would graph this line using the slope-intercept method.

OpenStudy (anonymous):

I'll help Is this algebra?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay lets walk through this

OpenStudy (anonymous):

All I need is the answer, there are 6 parts I will type the rest after this one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Y=-2/3x+400. (slope)m=-2/3 (Y-intercept)b=400 really simple

OpenStudy (anonymous):

So here is part one Change the equation into slope-intercept form. Identify the slope and y-intercept of the equation. Be sure to show all of your work.

OpenStudy (anonymous):

if you want to know how i got it well you have to take the equation 2x + 3y = 1,200 -2x---> 3y= -2x+1,200 now divide 3 by -2x and 1,200 you get---> y= -2/3x +400

OpenStudy (anonymous):

Ok here is the second part

OpenStudy (anonymous):

to find the slope and y-intercept use this y=mx+b (m is slope, b is y-intercept.) y= -2/3x(m) +400(b)

OpenStudy (anonymous):

Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

Well you would plot point (400,0) then use your slope -2/3 (rise over run theory) from point (400,0) go down 2 on the Y axis then to the right 3 points on the x axis. You should have cord (398,3) now draw the line by connecting the dots.

OpenStudy (anonymous):

Now part 3: Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences.